OpenStudy (anonymous):

in a graduating class having the same number of boys and girls the guidance counselor finds that 1/2 the boys and 1/3 of the girls are going to collage. what percent of the class is going to collage? A. 33.3% B. 41.7% C. 58.3% D. 66% E. 83.3%

OpenStudy (debbieg):

\[\frac{ 1 }{ 2 }\times\frac{ 1 }{ 2 }+\frac{ 1 }{ 3 }\times\frac{ 1 }{ 2 }=\frac{ 1 }{ 4 }+\frac{ 1 }{ 6 }=?\] Use LCD to find the sum, then convert to a percent.

OpenStudy (anonymous):

is it 41.7%?

OpenStudy (debbieg):

Yup! :)